Use the following corn futures quotes:  

  Corn 5,000 bushels
  Contract Month Open High Low Settle Chg Open Int
  Mar 455.125 457.000 451.750 452.000 −2.750 597,913  
  May 467.000 468.000 463.000 463.250 −2.750 137,547  
  July 477.000 477.500 472.500 473.000 −2.000 153,164  
  Sep 475.000 475.500 471.750 472.250 −2.000 29,258  

   

Suppose you buy 22 of the September corn futures contracts at the last price of the day. One month from now, the futures price of this contract is 463.25, and you close out your position. Calculate your dollar profit on this investment. (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Initial value of position = 22*(5,000)($4.7225) = $519,475.00

Final value of position = 22(5,000)($4.6325) = $509,575.00

Dollar profit = $509,575.00 – $519,475.00 = –$9,900.00